# GREEN: Generative Radiology Report Evaluation and Error Notation

### v1

@@ A.4 Algorithm for Modifying Radiology Reports
1def get_prompt(self, report):
2 error_types = self.get_error_combination(report)
3 # randomly choose if subtle or not, if subtle add "sentence"
4 subtle_change = ""
5 if random.random() > 0.5:
6 subtle_change = "Aim for subtlety, adjusting only one word where feasible. "
7 if not error_types == "no errors":
8 return f"[Objective]: Create a candidate radiology report that subtly integrates specific errors based on the provided reference report.
9 Process Overview: You will be presented with:
10 1. Style of errors.
11 2. A reference radiology report to base your candidate report on.
12 3. The desired format for your candidate report. Note: Be short in your response!
13
14 Style of errors:
15 Introduce errors related to {error_types}. The errors should be woven into the report as if they were genuine observations from a medical image, without any meta-commentary on their accuracy. {subtle_change}
16
@@ 14 Style of errors:
18
19 return f"[Objective]: Create a candidate radiology report that has the same
20 clinical meaning but is slightly rephrased.

### v2

@@ 7 Limitations
Acknowledgments
S.O. receives research support by the German Research Foundation ((ID: 517316550). A.C. receives research support from R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060, AY2AX000045, and contracts 75N92020C00008, 75N92020C00021. C.B. receives research support from Promedica Foundation, Switzerland.
This work was also supported in part by the Medical Imaging and Data Resource Center (MIDRC), which is funded by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) of the National Institutes of Health under contract 75N92020D00021 and through The Advanced Research Projects Agency for Health (ARPA-H).
@@ A.4 Algorithm for Modifying Radiology Reports
def get_prompt(self, report):
error_types = self.get_error_combination(report)
# randomly choose if subtle or not, if subtle add "sentence"
subtle_change = ""
if random.random() > 0.5:
subtle_change = "Aim for subtlety, adjusting only one word where feasible. "
if not error_types == "no errors":
return f"[Objective]: Create a candidate radiology report that subtly integrates specific errors based on the provided reference report.
Process Overview: You will be presented with:
1. Style of errors.
2. A reference radiology report to base your candidate report on.
3. The desired format for your candidate report. Note: Be short in your response!
Style of errors:
Introduce errors related to {error_types}. The errors should be woven into the report as if they were genuine observations from a medical image, without any meta-commentary on their accuracy. {subtle_change}
@@
return f"[Objective]: Create a candidate radiology report that has the same
clinical meaning but is slightly rephrased.
@@ A.5 Visualization of the GREEN Summary Clustering Technique
A.6 Test for difference in location of the mean errors
Table 10: Paired Wilcoxon test for a significant difference in error counts of experts to the mean expert
Expert W Statistic P-value
Expert 0 906.0 $1.83\times 10^{-18}$
Expert 1 1408.0 $4.77\times 10^{-14}$
Expert 2 2265.0 $6.22\times 10^{-9}$
Expert 3 2450.0 $2.01\times 10^{-8}$
Expert 4 3130.0 $4.76\times 10^{-5}$
Expert 5 1505.5 $2.11\times 10^{-13}$
Table 11: Paired Wilcoxon test for a significant difference in error counts of models to the mean expert
Model W Statistic P-value
Mistral-v0.1 (7B) 722.50 $4.30\times 10^{-26}$
LaMA-2 (7B) 2324.50 $3.41\times 10^{-19}$
Phi-2 (2.7B) 2294.50 $4.00\times 10^{-19}$
RadLLaMA-2 (7B) 2345.50 $2.19\times 10^{-9}$
RadPhi-2 (2.7B) 1728.00 $4.63\times 10^{-20}$
GREEN GPT-4 2554.00 $8.93\times 10^{-9}$
A.7 Fine-grained interexpert correlation